OpenStudy (anonymous):

The perimeter of a rectangle is 24 inches. Find the dimensions if its length is 3 inches greater than its width. a) The equation that best describes this is: b) The dimensions of the rectangle are:

OpenStudy (anonymous):

someone help!

OpenStudy (anonymous):

7,5&4,5

OpenStudy (anonymous):

formula: 2l +2w, then l= 3+w, substituting, 2(3+w)+2w=24 6+2w+2w=24 6+4w=24 4w=24-6 4w=18 w=18/4 or w=9/2, l=3+w l=3+9/2 l=15/2 gets?

OpenStudy (anonymous):

length=9 inches width=3 inches